A right triangle has a hypotenuse of 17 inches and a leg of 8 inches. What is the missing leg?
Answer by stanbon(75887)

A right triangle has a hypotenuse of 17 inches and a leg of 8 inches. What is the missing leg?
----
Let the missing leg have length x inches.
---
Solve:
x^2 + 8^2 = 17^2
----
x^2 = 225
----
x = 15 inches
==================
Cheers,
Stan H.
===========